In general, we take a – rd, a – (r – 1)d, …., a – d, a, a + rd in case we have to take (2r + 1) terms in an A.P. what does it mean
 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Madhur
 
T(2r + 1) = a + (2r + 1 – 1)d
 
 = a + 2r d
 
In case of any query please feel free to ask
 
Regards
Arun (askIITians forum expert)